# Resilience for Regular Path Queries: Towards a Complexity Classification

### v2

@@ 5.1. Four-Legged Languages
  • We have $\delta_{2}\neq\epsilon$. Then we claim that $\alpha\coloneq\gamma^{\prime}$, $\beta\coloneq\delta^{\prime}$, $\gamma\coloneq\alpha_{1}$ and $\delta\coloneq\delta_{1}$ are stable legs. Indeed, observe first that none of $\alpha,\beta,\gamma,\delta$ is empty. Moreover we have $\alpha x\beta=\gamma^{\prime}x\delta^{\prime}\in L$ and $\gamma x\delta=\alpha_{1}x\delta_{1}\in L$. Last, no infix of $\alpha x\delta=\gamma^{\prime}x\delta_{1}$ can be in $L$ because $\gamma^{\prime}x\delta_{1}\delta_{2}=\gamma^{\prime}x\delta^{\prime}$ is in $L$ and $\delta_{2}\neq\epsilon$ so any infix of $\alpha x\delta$ is a strict infix of a word of $L$; but $L$ is reduced so no strict infix of a word of $L$ is a word of $L$, hence no infix of $\alpha x\delta$ is a word of $L$.
  • We have $\alpha_{2}\neq\epsilon$. Then we do the symmetric choice of taking $\alpha\coloneq\alpha_{1},\beta\coloneq\delta_{1}$, $\gamma\coloneq\alpha^{\prime}$, and $\delta\coloneq\beta^{\prime}$ and claim that they are stable legs. Instead, none of $\alpha,\beta,\gamma,\delta$ is empty. Further, $\alpha x\beta=\alpha_{1}a\delta_{1}$ and $\gamma x\delta=\alpha^{\prime}x\beta^{\prime}$ are in $L$. Moreover we have $\alpha x\delta=\alpha_{1}x\beta^{\prime}$ and all its infixes are infixes of $\alpha^{\prime}x\beta^{\prime}$ which are strict because $\alpha_{2}\neq\epsilon$; as $\alpha^{\prime}x\beta^{\prime}\in L$ and $L$ is reduced, no infix of $\alpha x\delta$ is a word of $L$.
@@
For completeness, we point out that Section 5.2 does not hold if $L$ is not reduced:
{remark}
Take $L=\{x,axb,cxd\}$, which is four-legged, with the only possible body being $x$ and legs $(\alpha,\beta,\gamma,\delta)=(a,b,c,d)$; but these legs are not stable because $x$ is a infix of $\alpha x\delta=axd$ that is in the language.
Figure 6. Gadget for Case 1 of Section 5.2.
Figure 7. Gadget for Case 2 of Section 5.2.
We can now prove Section 5.2:
Proof of Section 5.2.
Let $L$ be a reduced and four-legged language. Using Section 5.2, let $x\in\Sigma$ be a body and $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}\in\Sigma^{+}$ be four stable legs, i.e., we have that $\alpha^{\prime}x\beta^{\prime}\in L$ and $\gamma^{\prime}x\delta^{\prime}\in L$ but no infix of $\alpha^{\prime}x\delta^{\prime}$ is in $L$. We then distinguish two cases:
@@
The first case is when, furthermore, no infix of $\gamma^{\prime}x\beta^{\prime}$ is in $L$. Then in this case, let us write $\alpha^{\prime}=a\alpha$ (i.e., $a$ is the first letter of $\alpha^{\prime}$, which is well defined since $\alpha^{\prime}\neq\epsilon$) and likewise write $\beta^{\prime}=\beta b$ and $\gamma^{\prime}=c\gamma$ and $\delta^{\prime}=\delta d$ (note that some of the letters $a$, $b$, $c$, and $d$ may be equal). In this case we claim that the gadget depicted in Figure 6 is a valid gadget. Edges in the gadgets labelled with words $\alpha,\beta,\gamma,\delta\in\Sigma^{*}$ represent chains labelled with the letters in the words. If any of the words is empty, then the head node of the edge is merged with the tail node. Such a drawing allows us to show hardness of any language that can be obtained by replacing the $\alpha,\beta,\gamma,\delta$ with any words in $\Sigma^{*}$. This is possible since, for any language that meets the conditions of case 1, a condensed hypergraph of matches can be obtained that does not contain any vertices corresponding to a word (i.e., a Greek symbols). We check this programmatically by exhaustively verifying that all directed paths in the completed gadget are either accounted for by the condensed hypergraph of matches or are not in the language. Thus, we know that for any such language, a condensed hypergraph of matches will be an odd chain as shown in Figure 6, and hence the gadget is a valid gadget.
@@
The second case is when there are some infixes of $\gamma^{\prime}x\beta^{\prime}$ that are in $L$. Observe that such infixes must intersect non-trivially both $\gamma^{\prime}$ and $\beta^{\prime}$. Indeed, infixes of $\gamma^{\prime}x\beta^{\prime}$ which do not intersect $\gamma^{\prime}$ are strict infixes of $\alpha^{\prime}x\beta^{\prime}\in L$ so they are not in $L$ because $L$ is reduced, and likewise infixes of $\gamma^{\prime}x\beta^{\prime}$ which do not intersect $\beta^{\prime}$ are strict infixes of $\gamma^{\prime}x\delta^{\prime}\in L$ so again they are not in $L$. Let us write $\alpha^{\prime}=a\alpha$ (i.e., $a$ is the first letter of $\alpha^{\prime}$) and $\beta^{\prime}=b\beta$ ($b$ is the first letter of $\beta^{\prime}$) and $\gamma^{\prime}=c_{1}\gamma_{1}$ ($c_{1}$ is the first letter of $\gamma^{\prime}$) and $\gamma^{\prime}=\gamma_{2}c_{2}$ ($c_{2}$ is the last letter of $\gamma^{\prime}$) and $\delta^{\prime}=\delta d$ ($d$ is the last letter of $\delta^{\prime}$). Then we claim that the gadget depicted in Figure 7 is a valid gadget. Indeed, we can programmatically verify this as well by exhaustively checking all directed paths. We observe that all paths (1) either contain or are subsumed by $\alpha^{\prime}x\beta^{\prime}$ or $\gamma^{\prime}x\delta^{\prime}$ matches, (2) are an infix of $\alpha^{\prime}x\delta^{\prime}$ (and hence can never be a match) or (3) contain $c_{2}xb$ as an infix. We notice that any infix of $\gamma^{\prime}a\beta^{\prime}$ that is a match must contain $c_{2}xb$ as an infix, and we know that such a match exists. For any such match, the other letters of the match can be eliminated in the condensed hypergraph of matches, as they are subsumed (via the edge domination rule) by the $c_{2}xb$ thus the node domination rule can remove the other letters of the $\gamma^{\prime}a\beta^{\prime}$-infix match. Thus, it suffices to consider $\alpha^{\prime}x\beta^{\prime}$, $\gamma^{\prime}x\delta^{\prime}$, and $c_{2}xb$ as the only matches in the condensed hypergraph of matches. When we do so, we can see that there exists a condensed hypergraph of matches that is an odd chain as shown in Figure 7, and hence the gadget is valid.
This concludes the proof of Section 5.2. ∎
Note that Section 5.2 is not restricted to regular languages and applies to arbitrary languages (even when resilience is not in NP). Next, we spell out two consequences of the result.
@@ 5.3. Implications of Hardness for Four-Legged Languages
{toappendix}
5.4. Reduced Languages of Star-Free Languages are Star-Free
We formally show the following result stated in the main text:
{claim}
Let $L$ be a star-free regular language. Then $\mathtt{reduce}(L)$ is star-free.
Of course, the converse statement is not true, e.g., $(aa)^{*}$ is not star-free but its reduced language $\epsilon$ is star-free. Let us prove the result:
Proof of Section 5.4.
We use an alternative characterization of star-free languages (mpri): they are those languages that can be constructed inductively by applying the operations of concatenation and the Boolean operations (including complementation) on base languages consisting of the empty language, the singleton language containing only the empty word, and the singleton languages formed of single-letter words for each letter of the alphabet $\Sigma$.
We then use the fact that, for any language $L$, we can express $\mathtt{reduce}(L)$ as follows:
$\mathtt{reduce}(L)=L\setminus(\Sigma^{+}L\Sigma^{*}\cup\Sigma^{*}L\Sigma^{+})$
Note that $\Sigma^{*}$ is the complement of the empty set, and $\Sigma^{+}$ is the concatenation of $\Sigma^{*}$ with the union of all letters of $\Sigma$, so both are star-free languages. Hence, if $L$ is a star-free language, then thanks to the fact that star-free languages are closed under concatenation and Boolean operations, so is $\mathtt{reduce}(L)$, which concludes the proof. ∎
5.5. Proof of Section 5.5: Hardness of Resilience for Non-Star-Free Languages
@@ 5.4. Reduced Languages of Star-Free Languages are Star-Free
We say that a regular language $L$ is star-free if there exists a $k>0$ such that, for every $\rho,\sigma,\tau\in\Sigma^{*}$, for every $m\geq k$, we have $\rho\sigma^{k}\tau\in L$ iff $\rho\sigma^{m}\tau\in L$. Star-free languages have many equivalent characterizations (see, e.g., (mpri)), and if a language $L$ is star-free then so is $\mathtt{reduce}(L)$ (see Section 5.4). We can show that being star-free is a necessary condition to enjoy tractable resilience, because non-star-free reduced regular languages are four-legged and hence resilience for them is NP-hard by Section 5.2:
{lemmarep}
[Hardness of non-star-free languages] Let $L$ be a regular language which is reduced and non-star-free. Then $L$ is four-legged.
@@
We use the equivalent characterization of star-free languages as those languages that are counter-free (mcnaughton1971counter). Let $A$ be a DFA which recognizes $L$, and let $k^{\prime}>0$ be greater than the number of states of $A$. Let $\rho,\sigma,\tau\in\Sigma^{*}$ and $m\geq k$ be witnesses to the fact that $L$ is not counter-free for $k=k^{\prime}$, i.e., exactly one of $\rho\sigma^{k}\tau$ and $\rho\sigma^{m}\tau$ is in the language. In particular this implies that $\sigma\neq\epsilon$; let us write $\sigma=x\sigma^{\prime}$.
By the pigeonhole principle, there must be $0\leq i<j\leq k^{\prime}$ such that the partial runs of $\rho\sigma^{i}$ and $\rho\sigma^{j}$ finish at the same state when read by $A$. Let $\omega\coloneq j-i$. We know that we have $\rho\sigma^{k}\tau\in L$ iff $\rho\sigma^{k+\omega}\tau\in L$, because exchanging the prefixes $\rho\sigma^{i}$ and $\rho\sigma^{j}$ does not change the state that we reach. Iterating the argument, we get the following fact, dubbed (*): for any $p\geq 0$ we have $\rho\sigma^{k}\tau\in L$ iff $\rho\sigma^{k+p\omega}\tau\in L$. For similar reasons, we get the fact dubbed (**): for any $p\geq 0$ we have $\rho\sigma^{m}\tau\in L$ iff $\rho\sigma^{m+p\omega}\tau\in L$.
@@
  • $\beta=\sigma^{\prime}\sigma^{\phi}w$,
@@
  • $\delta=\sigma^{\prime}\sigma^{\psi}w$.
@@
  • $\alpha x\delta\in L$ because $\rho\sigma^{\psi}\tau\notin L$
This establishes that $L$ is four-legged. As it is reduced, we can conclude by Proposition 5.2. ∎
{toappendix}
5.6. Proof of Section 5.6: Dichotomy for Languages with Neutral Letters
@@
A second consequence of Section 5.2 concerns languages $L$ with a neutral letter, i.e., a letter $e\in\Sigma$ such that inserting or deleting $e$ anywhere does not change the membership of words to $L$. Formally, $e$ is neutral for $L$ if for every $\alpha,\beta\in\Sigma^{*}$ we have $\alpha\beta\in L$ iff $\alpha e\beta\in L$. In formal language theory, one often assumes the existence of a neutral letter as a technical assumption on languages (koucky2005bounded; barrington2005first). Under this assumption, we can show:
{propositionrep}
[Dichotomy for languages with neutral letters] Let $L$ be a language with a neutral letter. If $\mathtt{reduce}(L)$ is local then $\mathtt{RES}_{\mathrm{bag}}(L)$ is PTIME, otherwise $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard.
{proofsketch}
If $\mathtt{reduce}(L)$ is local then we have tractability thanks to Theorem 3.3. Otherwise, we can show that $\mathtt{reduce}(L)$ is either four-legged or contains a word of the form $aa$ (or both), so that at least one of Proposition 4.1 and Section 5.2 applies.
{toappendix}
To prove this result, we first show the claim mentioned in the proof sketch: for languages featuring a neutral letter, whenever the reduced language is non-local then it must either be four-legged or contain a word with a repeated letter, specifically a word of the form $aa$ for $a\in\Sigma$:
{lemmarep}
Let $L$ be a language with a neutral letter such that $\mathtt{reduce}(L)$ is not local. Then one of the following two cases holds (possibly both):
  • $\mathtt{reduce}(L)$ is four-legged;
  • there is $x\in\Sigma$ such that $xx\in\mathtt{reduce}(L)$.
To prove this result, let us prove the following intermediate result:
{claim}
Let $L$ be a language with a neutral letter $e$ and let $L^{\prime}=\mathtt{reduce}(L)$. Then for any nonempty words $\theta,\eta\in\Sigma^{+}$, if $\theta\eta\in L^{\prime}$ then $\theta e\eta\in L^{\prime}$.
@@
  • $\tau$ is an infix of $\theta$. Then it is a strict infix of $\theta\eta\in L^{\prime}$, which contradicts the fact that $L^{\prime}$ is reduced, so this case is impossible.
@@
Now, we know that $\tau$ is a strict infix of $\theta e\eta$ which includes the $e$. Let $\tau^{\prime}$ be obtained from $\tau$ by removing the $e$. We have $\tau^{\prime}\in L$ because $\tau\in L$ and $e$ is neutral for $L$. What is more, it is easily seen that $\tau^{\prime}$ is an infix of $\theta\eta$, and it is a strict infix because $\tau$ includes the $e$ so from $|\tau|<|\theta e\eta|$ (because $\tau$ is a strict infix) we deduce $|\tau^{\prime}|=|\tau|-1<|\theta e\eta|-1=|\theta\eta|$. So $\tau^{\prime}\in L$ is a strict infix of $\theta\eta\in L^{\prime}$, which contradicts the assumption that $L^{\prime}$ is reduced. ∎
We can now prove Section 5.6:
Proof of Section 5.6.
@@
Now, because $L^{\prime}$ is reduced, it is easy to check that we must have the following claim, dubbed ($\star$): we have $\alpha^{\prime}\beta^{\prime}\neq\epsilon$ and $\gamma^{\prime}\delta^{\prime}\neq\epsilon$ and $\alpha^{\prime}\gamma^{\prime}\neq\epsilon$ and $\beta^{\prime}\delta^{\prime}\neq\epsilon$. Indeed:
  • If $\alpha^{\prime}\beta^{\prime}=\epsilon$ then $\alpha^{\prime}x\beta^{\prime}\in L^{\prime}$ is equal to $x$ which is an infix of $\gamma^{\prime}x\delta^{\prime}\in L^{\prime}$, so by reducedness it must be the case that $\gamma^{\prime}\delta^{\prime}=\epsilon$, but now $x=\alpha^{\prime}x\delta^{\prime}=x$ and we assumed $x=\alpha^{\prime}x\delta^{\prime}\notin L^{\prime}$, contradiction.
  • If $\gamma^{\prime}\delta^{\prime}=\epsilon$, the reasoning is similar: in this case $\gamma^{\prime}x\delta^{\prime}$ is equal to $x$, which is an infix of $\alpha^{\prime}x\beta^{\prime}$, so by reducedness we must have $\alpha^{\prime}\beta^{\prime}=\epsilon$, but then $x=\alpha^{\prime}x\beta^{\prime}\in L^{\prime}$ and $x=\alpha^{\prime}x\delta^{\prime}\notin L^{\prime}$ is a contradiction.

### v3

@@ 5.1. Four-Legged Languages
  • We have $\delta_{2}\neq\epsilon$. Then we claim that $\alpha\coloneq\gamma^{\prime}$, $\beta\coloneq\delta^{\prime}$, $\gamma\coloneq\alpha_{1}$ and $\delta\coloneq\delta_{1}$ are stable legs. Indeed, observe first that none of $\alpha,\beta,\gamma,\delta$ is empty. Moreover we have $\alpha x\beta=\gamma^{\prime}x\delta^{\prime}\in L$ and $\gamma x\delta=\alpha_{1}x\delta_{1}\in L$. Last, no infix of $\alpha x\delta=\gamma^{\prime}x\delta_{1}$ can be in $L$ because $\gamma^{\prime}x\delta_{1}\delta_{2}=\gamma^{\prime}x\delta^{\prime}$ is in $L$ and $\delta_{2}\neq\epsilon$ so any infix of $\alpha x\delta$ is a strict infix of a word of $L$; but $L$ is infix-free so no strict infix of a word of $L$ is a word of $L$, hence no infix of $\alpha x\delta$ is a word of $L$.
  • We have $\alpha_{2}\neq\epsilon$. Then we do the symmetric choice of taking $\alpha\coloneq\alpha_{1},\beta\coloneq\delta_{1}$, $\gamma\coloneq\alpha^{\prime}$, and $\delta\coloneq\beta^{\prime}$ and claim that they are stable legs. Instead, none of $\alpha,\beta,\gamma,\delta$ is empty. Further, $\alpha x\beta=\alpha_{1}x\delta_{1}$ and $\gamma x\delta=\alpha^{\prime}x\beta^{\prime}$ are in $L$. Moreover we have $\alpha x\delta=\alpha_{1}x\beta^{\prime}$ and all its infixes are infixes of $\alpha^{\prime}x\beta^{\prime}$ which are strict because $\alpha_{2}\neq\epsilon$; as $\alpha^{\prime}x\beta^{\prime}\in L$ and $L$ is infix-free, no infix of $\alpha x\delta$ is a word of $L$.
@@
Figure 5. Completed gadget for Case 1 of Theorem 5.3.
Figure 6. Completed gadget for Case 2 of Theorem 5.3.
We can now prove that resilience is hard for four-legged languages (Theorem 5.3):
Proof of Theorem 5.3.
Let $L$ be an infix-free and four-legged language. Using Lemma 5.5, let $x\in\Sigma$ be a body and $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}\in\Sigma^{+}$ be four stable legs, i.e., we have that $\alpha^{\prime}x\beta^{\prime}\in L$ and $\gamma^{\prime}x\delta^{\prime}\in L$ but no infix of $\alpha^{\prime}x\delta^{\prime}$ is in $L$. We then distinguish two cases:
@@
The first case is when, furthermore, no infix of $\gamma^{\prime}x\beta^{\prime}$ is in $L$. Then in this case, let us write $\alpha^{\prime}=a\alpha$ (i.e., $a$ is the first letter of $\alpha^{\prime}$, which is well defined since $\alpha^{\prime}\neq\epsilon$) and likewise write $\beta^{\prime}=\beta b$ and $\gamma^{\prime}=c\gamma$ and $\delta^{\prime}=\delta d$ (note that some of the letters $a$, $b$, $c$, and $d$ may be equal). In this case we claim that the (completed) gadget depicted in Figure 5 is a valid gadget. [^5] We point out that this gadget is essentially the generalization of the gadget in Figure 4(a) for $L=axb|cxd$. Edges in the gadgets labeled with words $\alpha,\beta,\gamma,\delta\in\Sigma^{*}$ represent paths labeled with the letters in the words. If any of the words is empty, then the head node of the edge is merged with the tail node. Such a drawing allows us to show hardness of any language that can be obtained by replacing the $\alpha,\beta,\gamma,\delta$ with any words in $\Sigma^{*}$. This is possible since, for any language that meets the conditions of case 1, a condensed hypergraph of matches can be obtained that does not contain any vertices corresponding to a word (i.e., a Greek symbol). We can then check exhaustively that all directed walks in the completed gadget are either accounted for by the condensed hypergraph of matches or are not in the language. Thus, we know that for any such language, a condensed hypergraph of matches will be an odd path as shown in Figure 5, and hence the gadget is a valid gadget.
@@
The second case is when there are some infixes of $\gamma^{\prime}x\beta^{\prime}$ that are in $L$. Observe that such infixes must intersect non-trivially both $\gamma^{\prime}$ and $\beta^{\prime}$. Indeed, infixes of $\gamma^{\prime}x\beta^{\prime}$ which do not intersect $\gamma^{\prime}$ are strict infixes of $\alpha^{\prime}x\beta^{\prime}\in L$ so they are not in $L$ because $L$ is infix-free, and likewise infixes of $\gamma^{\prime}x\beta^{\prime}$ which do not intersect $\beta^{\prime}$ are strict infixes of $\gamma^{\prime}x\delta^{\prime}\in L$ so again they are not in $L$. Let us write $\alpha^{\prime}=a\alpha$ (i.e., $a$ is the first letter of $\alpha^{\prime}$) and $\beta^{\prime}=b\beta$ ($b$ is the first letter of $\beta^{\prime}$) and $\gamma^{\prime}=c_{1}\gamma_{1}$ ($c_{1}$ is the first letter of $\gamma^{\prime}$) and $\gamma^{\prime}=\gamma_{2}c_{2}$ ($c_{2}$ is the last letter of $\gamma^{\prime}$) and $\delta^{\prime}=\delta d$ ($d$ is the last letter of $\delta^{\prime}$). Then we claim that the (completed) gadget depicted in Figure 6 is a valid gadget. Indeed, we can again check exhaustively all directed walks in the gadget and observe that the following holds for all of them: they (1) either contain or are subsumed by $\alpha^{\prime}x\beta^{\prime}$ or $\gamma^{\prime}x\delta^{\prime}$ matches; (2) are an infix of $\alpha^{\prime}x\delta^{\prime}$ (and hence can never be a match); or (3) contain $c_{2}xb$ as an infix. We notice that any infix of $\gamma^{\prime}x\beta^{\prime}$ that is a match must contain $c_{2}xb$ as an infix, and we know that such a match exists. For any such match, the other letters of the match can be eliminated in the condensed hypergraph of matches, as they are subsumed (via the edge domination rule) by the $c_{2}xb$ - thus the node domination rule can remove the other letters of the $\gamma^{\prime}x\beta^{\prime}$-infix match. Thus, it suffices to consider $\alpha^{\prime}x\beta^{\prime}$, $\gamma^{\prime}x\delta^{\prime}$, and $c_{2}xb$ as the only matches in the condensed hypergraph of matches. When we do so, we can see that there exists a condensed hypergraph of matches that is an odd path as shown in Figure 6, and hence the gadget is valid.
This concludes the proof of Theorem 5.3. ∎
Note that Theorem 5.3 is not restricted to regular languages and applies to arbitrary languages (even when resilience is not in NP).
@@ 5.2. Implications of Hardness for Four-Legged Languages
Next, we spell out two consequences of Theorem 5.3. The first one concerns non-star-free languages, the second concerns languages with a neutral letter.
@@
We say that a regular language $L$ is star-free if there exists a $k>0$ such that, for every $\rho,\sigma,\tau\in\Sigma^{*}$, for every $m\geq k$, we have $\rho\sigma^{k}\tau\in L$ iff $\rho\sigma^{m}\tau\in L$. Star-free languages have many equivalent characterizations (see, e.g., (mpri, )), including being counter-free (mcnaughton1971counter, ). Further, if a language $L$ is star-free then so is $\mathtt{IF}(L)$ (see LABEL:apx:starfreered). We can show that being star-free is a necessary condition to enjoy tractable resilience, because non-star-free infix-free regular languages are four-legged and hence resilience for them is NP-hard by Theorem 5.3:
Lemma 5.6 (Hardness of non-star-free languages).
Let $L$ be a regular language which is infix-free and non-star-free. Then $L$ is four-legged.
@@
Let $A$ be a DFA which recognizes $L$, and let $k>0$ be greater than the number of states of $A$. From the fact that $L$ is not star-free, applying the definition to this choice of $k$, we obtain $\rho,\sigma,\tau\in\Sigma^{*}$ and $m\geq k$ such that exactly one of $\rho\sigma^{k}\tau$ and $\rho\sigma^{m}\tau$ is in the language. In particular this implies that $\sigma\neq\epsilon$; let us write $\sigma=x\sigma^{\prime}$.
As $k$ is greater than the number of states of $A$, by the pigeonhole principle, there must be $0\leq i<j\leq k$ such that the partial runs of $\rho\sigma^{i}$ and $\rho\sigma^{j}$ finish at the same state when read by $A$. Let $\omega\coloneq j-i$. We know that we have $\rho\sigma^{k}\tau\in L$ iff $\rho\sigma^{k+\omega}\tau\in L$, because exchanging the prefixes $\rho\sigma^{i}$ and $\rho\sigma^{j}$ does not change the state that we reach. Iterating the argument, we get the following fact, dubbed (*): for any $p\geq 0$ we have $\rho\sigma^{k}\tau\in L$ iff $\rho\sigma^{k+p\omega}\tau\in L$. For similar reasons, we get the fact dubbed (**): for any $p\geq 0$ we have $\rho\sigma^{m}\tau\in L$ iff $\rho\sigma^{m+p\omega}\tau\in L$.
@@
  • $\beta=\sigma^{\prime}\sigma^{\phi}\tau$,
@@
  • $\delta=\sigma^{\prime}\sigma^{\psi}\tau$.
@@
  • $\alpha x\delta\notin L$ because $\rho\sigma^{\psi}\tau\notin L$
This establishes that $L$ is four-legged. As it is infix-free, we can conclude by Proposition 5.3. ∎
@@
A second consequence of Theorem 5.3 concerns languages $L$ with a neutral letter, i.e., a letter $e\in\Sigma$ such that inserting or deleting $e$ anywhere does not change the membership of words to $L$. Formally, $e$ is neutral for $L$ if for every $\alpha,\beta\in\Sigma^{*}$ we have $\alpha\beta\in L$ iff $\alpha e\beta\in L$. In formal language theory, one often assumes the existence of a neutral letter as a technical assumption on languages (koucky2005bounded, ; barrington2005first, ). Under this assumption, we can show:
Proposition 5.7 (Dichotomy for languages with neutral letters).
Let $L$ be a language with a neutral letter. If $\mathtt{IF}(L)$ is local then $\mathtt{RES}_{\mathrm{bag}}(L)$ is PTIME, otherwise $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard.
Note that we can read this result as implying that, in the presence of a neutral letter, there are not many tractable cases left. Indeed, intuitively, when considering a language $L$ with neutral letter whose infix-free sublanguage is local, then membership to $L$ must be guided only by the first and last non-neutral letters and the set of non-neutral letters that occur in the word.
We prove Proposition 5.7 in the remaining of Section 5.2. To prove this result, we first show that, for languages featuring a neutral letter, whenever the infix-free language is non-local then it must either be four-legged or contain a word with a repeated letter, specifically a word of the form $aa$ for $a\in\Sigma$:
Lemma 5.8.
Let $L$ be a language with a neutral letter such that $\mathtt{IF}(L)$ is not local. Then one of the following two cases holds (possibly both):
  • $\mathtt{IF}(L)$ is four-legged;
  • there is $x\in\Sigma$ such that $xx\in\mathtt{IF}(L)$.
To prove this lemma, let us prove the following intermediate result:
Claim 5.9.
Let $L$ be a language with a neutral letter $e$ and let $L^{\prime}=\mathtt{IF}(L)$. Then for any non-empty words $\theta,\eta\in\Sigma^{+}$, if $\theta\eta\in L^{\prime}$ then $\theta e\eta\in L^{\prime}$.
@@
  • $\tau$ is an infix of $\theta$. Then it is a strict infix of $\theta\eta\in L^{\prime}$, which contradicts the fact that $L^{\prime}$ is infix-free, so this case is impossible.
@@
Now, we know that $\tau$ is a strict infix of $\theta e\eta$ which includes the $e$. Let $\tau^{\prime}$ be obtained from $\tau$ by removing the $e$. We have $\tau^{\prime}\in L$ because $\tau\in L$ and $e$ is neutral for $L$. What is more, it is easily seen that $\tau^{\prime}$ is an infix of $\theta\eta$, and it is a strict infix because $\tau$ includes the $e$ so from $|\tau|<|\theta e\eta|$ (because $\tau$ is a strict infix) we deduce $|\tau^{\prime}|=|\tau|-1<|\theta e\eta|-1=|\theta\eta|$. So $\tau^{\prime}\in L$ is a strict infix of $\theta\eta\in L^{\prime}$, which contradicts the assumption that $L^{\prime}$ is infix-free. ∎
We can now prove Lemma 5.8:
Proof of Lemma 5.8.
@@
Now, because $L^{\prime}$ is infix-free, it is easy to check that we must have the following claim, dubbed ($\star$): we have $\alpha^{\prime}\beta^{\prime}\neq\epsilon$ and $\gamma^{\prime}\delta^{\prime}\neq\epsilon$ and $\alpha^{\prime}\gamma^{\prime}\neq\epsilon$ and $\beta^{\prime}\delta^{\prime}\neq\epsilon$. Indeed:
  • If $\alpha^{\prime}\beta^{\prime}=\epsilon$ then $\alpha^{\prime}x\beta^{\prime}\in L^{\prime}$ is equal to $x$ which is an infix of $\gamma^{\prime}x\delta^{\prime}\in L^{\prime}$, so by infix-freeness it must be the case that $\gamma^{\prime}\delta^{\prime}=\epsilon$, but now $x=\alpha^{\prime}x\delta^{\prime}=x$ and we assumed $x=\alpha^{\prime}x\delta^{\prime}\notin L^{\prime}$, contradiction.
  • If $\gamma^{\prime}\delta^{\prime}=\epsilon$, the reasoning is similar: in this case $\gamma^{\prime}x\delta^{\prime}$ is equal to $x$, which is an infix of $\alpha^{\prime}x\beta^{\prime}$, so by infix-freeness we must have $\alpha^{\prime}\beta^{\prime}=\epsilon$, but then $x=\alpha^{\prime}x\beta^{\prime}\in L^{\prime}$ and $x=\alpha^{\prime}x\delta^{\prime}\notin L^{\prime}$ is a contradiction.